OpenStudy (anonymous):

Can someone help me with this one? lim->-x⁻ 14x/x²-1

OpenStudy (anonymous):

ok what u do is multiply the numbers 

OpenStudy (anonymous):

\[\lim_{x \rightarrow -1⁻}\frac{ 14x }{ x²-1 }\]

OpenStudy (harsimran_hs4):

you can say that limit does not exit in this case....

OpenStudy (anonymous):

why?

OpenStudy (harsimran_hs4):

\[\frac{ 14x }{ (x+1) (x-1) } = \frac{ 14x }{ x-1 } \frac{ 1 }{ x+1 }\] with the 1st part it`s fine but our concentration is on 1/x+1 part so when x approaches -1 from left side 1/x+1 tends to -ve infinity and when x approaches -1 from right side 1/x+1 approaches +ve infinity \so LHL is not equal to RHL so limit does not exit or limit is not defined

OpenStudy (anonymous):

Ok that helps, thanks. I don't fully understand the relationship between these two new functions ant their composite function though and how I could find out that the limit from the negative side is not defined at -1, (when at -2 it would be -ve infinity)

OpenStudy (harsimran_hs4):

at -2 it won`t be -ve infinity rather i had just written the same function mid it it`s not the composite functions it was just done because i knew that left part would create no trouble in finding limit but only right one is the one we need to analyze...

OpenStudy (anonymous):

Ok thanks for help. I'm gonna close this and move on

OpenStudy (harsimran_hs4):

cool just think over it again some other time you will get hold of it